solve the equation
1/8(x-6)^3/2=1
please help me solve this equation

Answer by edjones(8007)

1/8(x-6)^(3/2)=1
(x-6)^(3/2)=8
sqrt(x-6)=2 cube root of each side.
x-6=4 square each side.
x=10
.
Ed